OpenStudy (anonymous):

Part 1: [4 points] Find the measures of central tendency (mean, median, mode and range) for the following data set: 9, 13, 10, 7, 16, 15, 11 Part 2: [3 points] Tell which of these measures would best represent the whole data set and why.

OpenStudy (anonymous):

Mean : Add up all the numbers and divide that by the amount of numbers in the data set. ie. (9 + 13 + 10 + 7 + 16 + 15 + 11) / 7 = 81/7 Median: First, sort the list from lowest to highest. Then, you know there is an odd amount of terms, so the median is the middle term, which is 11. 7, 9, 10 ||| 11|||| 12, 15, 16 Mode : the number that appears the most. In this set they all appear once. Range = max - min -> 16-7